Filter: |
Resolved
Get updates: RSS | Twitter

VMs Down/Restarted

Issue status: Resolved

Resolved at: 23 Jul 2015, 11:46 UTC

London peering offline

Issue status: Resolved

Resolved at: 5 Jul 2015, 08:59 UTC

A denial of service attack is affecting some connectivity.

Issue status: Resolved

Resolved at: 3 Jul 2015, 23:57 UTC

BigV head18 reboot (~ 1735 28 June 2015 UTC+1)

Issue status: Resolved

Resolved at: 29 Jun 2015, 14:22 UTC

Fasthosts DNS (livedns.co.uk) is broken

Issue status: Resolved

Resolved at: 18 Jun 2015, 19:02 UTC

Packet Loss

Issue status: Resolved

Resolved at: 18 Jun 2015, 18:02 UTC

Instability on BigV head 25

Issue status: Resolved

Resolved at: 14 Jun 2015, 22:06 UTC

Careers site unavailable

Issue status: Resolved

Resolved at: 3 Jun 2015, 17:06 UTC

IX Manchester port upgrade

Issue status: Resolved

Resolved at: 2 Jun 2015, 14:47 UTC

Issue still not addressed? Please contact support.